OpenStudy (anonymous):

Use linear approximation, i.e. the tangent line, to approximate cube root of 0.8 as follows: The equation of the tangent line to f (x) at x = 1 can be written in the form y =___? Using this, we find our approximation for cube root of 0.8 is = ____?

OpenStudy (anonymous):

Can you write tangent line equation?

OpenStudy (anonymous):

f(x ) = ³√ x -> f (1) = 1 f '(x) = 1/3 x ^(2/3) => f '(1) = 1/3 Tangent line: y = 1/3 ( x - 1) + 1 = x/3 + 2/3 Thus f ( .8 ) = 1/3 ( .8 + 2) = 2.8/ 3 = .7

OpenStudy (anonymous):

Oops, = .933333

OpenStudy (anonymous):

Pls give the feedback, so at least we know the result of our work!

OpenStudy (anonymous):

@Chlorophyll is right. I dont know why but i took the derivative of\[1/\sqrt[3]{x}\] :)

OpenStudy (anonymous):

@myko Our mind just trick us somehow! That's why I always need feed back :)

OpenStudy (anonymous):

Yeah...the first answer was correct, but not the second....I'm trying to see why it is .933333

OpenStudy (anonymous):

So far thanks for the help chlorophyll!!

OpenStudy (anonymous):

@Unam, I often have issue with calculation! That's why I asked your feedback about the previous one!

OpenStudy (anonymous):

Thanks..you got it right Chlorophyll...the computer needs is too picky about digits! Thanks so much!

OpenStudy (anonymous):

@Unam, make sure you give out the medal to confirm that our work is correct!

OpenStudy (anonymous):

I did press the "best answer," is that what you meant?